Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen: Overview, Questions, Preparation

Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen 2025 ( Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen )

Payal Gupta
Updated on Sep 8, 2025 14:16 IST

By Payal Gupta, Retainer

Table of contents
  • Probability Short Answer Type Questions
  • Probability Long Answer Type Questions
  • Probability Objective Type Questions
  • Probability True or False Type Questions
  • Probability Fill in the blank Type Questions
  • Probability Matching Type Questions
Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen Logo

Probability Short Answer Type Questions

1. If the letters of the word ALGORITHM are arranged at random in a row, what is the probability that the letters GOR must remain together as a unit?

Sol. W o r d     A L G O R I T H M     h a s     9     l e t t e r s . I f     G O R     r e m a i n     t o g e t h e r ,     t h e n     i t     w i l l     r e m a i n     t o g e t h e r . ∴     N u m b e r     o f     l e t t e r s = A L G O R I T H M = 6 + 1 = 7 N u m b e r     o f     w o r d s = 7 ! a n d     t h e     t o t a l     n u m b e r     o f     w o r d s     f r o m     A L G O R I T H M = 9 ! S o ,     t h e     r e q u i r e d     p r o b a b i l i t y = 7 ! 9 ! = 7 ! 9 . 8 . 7 ! = 1 7 2 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y = 1 7 2 .

2. Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randomly, what is the probability that the married couple will have nonadjacent desks?

[Hint: First find the probability that the couple has adjacent desks, and then subtract it from 1.]

Sol. W o r d     A L G O R I T H M     h a s     9     l e t t e r s . N u m b e r     o f     d e s k     o c c u p i e d     b y     o n e     c o u p l e = 1 O n l y ( 4 + 1 ) = 5     p e r s o n s     t o     b e     a s s i g n e d . ∴     N u m b e r     o f     w a y s     o f     a s s i g n i n g     t h e s e     5     p e r s o n s = 5 !   × 2 ! T o t a l     n u m b e r     o f     w a y s     o f     a s s i g n i n g     t h e s e     6     p e r s o n s = 6 ! ∴     Probability     t h a t     a     c o u p l e     h a s     a d j a c e n t     d e s k = 5 !   × 2 ! 6 ! = 1 3 S o ,     t h e     p r o b a b i l i t y     t h a t     t h e     m a r r i e d     c o u p l e     w i l l     h a v e     n o - a d j a c e n t     d e s k s = 1 − 1 3 = 2 3 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y = 2 3 .
Q&A Icon
Commonly asked questions
Q:  

Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randomly, what is the probability that the married couple will have nonadjacent desks?

[Hint: First find the probability that the couple has adjacent desks, and then subtract it from 1.]

Read more
A: 

This is a short answer type question as classified in NCERT ExemplarWord  ALGORITHM  has  9  letters.Number  of  desk  occupied  by  one  couple=1Only(4+1)=5  persons  to  be  assigned.∴  Number  of  ways  of  assigning  these  5  persons=5! ×2!Total  number  of  ways  of  assigning  these  6  persons=6!∴ Probability   that  a  couple  has  adjacent  desk=5! ×2!6!=13So,  the  probability  that  the  married  couple  will  have  noadjacent  desks=1−13=23Hence,  the  required  probability=23.

Q:  

If the letters of the word ALGORITHM are arranged at random in a row, what is the probability that the letters GOR must remain together as a unit?

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

W o r d     A L G O R I T H M     h a s     9     l e t t e r s . I f     G O R     r e m a i n     t o g e t h e r ,     t h e n     i t     w i l l     r e m a i n     t o g e t h e r . ∴     N u m b e r     o f     l e t t e r s = A L G O R I T H M = 6 + 1 = 7 N u m b e r     o f     w o r d s = 7 ! a n d     t h e     t o t a l     n u m b e r     o f     w o r d s     f r o m     A L G O R I T H M = 9 ! S o ,     t h e     r e q u i r e d     p r o b a b i l i t y = 7 ! 9 ! = 7 ! 9 . 8 . 7 ! = 1 7 2 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y = 1 7 2 .

Q:  

Suppose an integer from 1 through 1000 is chosen at random. Find the probability that the integer is a multiple of 2 or a multiple of 9.

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

W e     h a v e     m u l t i p l e s     o f     2 ,     f r o m     1     t o     1 0 0 0     a r e     2 , 4 , 6 , 8 , … , 1 0 0 0 L e t     n     b e     t h e     n u m b e r     o f     t e r m s H e r e     a = 2 ,     d = 2 ,     a n = 1 0 0 0                                                         a n = a + ( n − 1 ) d                                             1 0 0 0 = 2 + ( n − 1 ) 2 ⇒                                   1 0 0 0 = 2 n               ⇒ n = 5 0 0 N o w     m u l t i p l e s     o f     9 ,     f r o m     1     t o     1 0 0 0     a r e     9 , 1 8 , 2 7 , … , 9 9 9 H e r e     a = 9 ,     d = 9 ,     a m = 9 9 9                                 [ m     b e     t h e     n u m b e r     o f     t e r m s ]                                                         a m = a + ( m − 1 ) d ⇒                                         9 9 9 = 9 + ( m − 1 ) 9 ⇒                                         9 9 9 = 9 m                   ⇒ m = 1 1 1 N o w     m u l t i p l e s     o f     2     a n d     9     a r e     1 8 , 3 6 , 5 4 , … , 9 9 0 H e r e     a = 1 8 ,     d = 1 8 ,     a p = 9 9 9                                 [ p     b e     t h e     n u m b e r     o f     t e r m s ] ∴                                                 a p = a + ( p − 1 ) d ⇒                                       9 9 0 = 1 8 + ( p − 1 ) 1 8 ⇒                                       9 9 0 = 1 8 p                 ⇒ p = 5 5 ∴     N u m b e r     o f     m u l t i p l e s     o f     2     o r     9 = 5 0 0 + 1 1 1 − 5 5 = 5 5 6 ∴     R e q u i r e d     p r o b a b i l i t y = P ( E ) n ( E ) = 5 5 6 1 0 0 0 = 0 . 5 5 6 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y = 0 . 5 5 6

Q:  

An experiment consists of rolling a die until a 2 appears.

(i) How many elements of the sample space correspond to the event that the 2 appears on the k t h roll of the die?

(ii) How many elements of the sample space correspond to the event that the 2 appears not later than the k t h roll of the die?

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

N u m b e r     o f     S a m p l e     s p a c e = 6 ( i ) G i v e n     t h a t     2     a p p e a r s     o n     t h e     k t h     r o l l     o f     t h e     d i e .           S o     f i r s t     ( k − 1 ) t h     r o l l     h a v e     5     o u t     c o m e s     e a c h     a n d     k t h     r o l l     r e s u l t s     2     i . e .     o n l y     1     o u t c o m e . ∴     N u m b e r     o f     e l e m e n t     o f     s a m p l e     s p a c e     c o r r e s p o n d     t o     t h e     e v e n t     t h a t     2     a p p e a r s     o n     t h e     k t h           r o l l     o f     t h e     d i e = 5 k − 1 . ( i i ) I n     t h i s     c a s e ,     2     a p p e a r s     n o t     l a t e r     t h a n     k t h     r o l l     o f     t h e     d i e ,     t h e n     i t     i s     p o s s i b l e     t h a t     2     c o m e s        in  first  throw  then  outcomes  will  be  5  and  2  outcomes  in  second  throw  i.e.  1  outcome. ∴         P o s s i b l e     o u t c o m e = 5 × 1 = 5       Similarly,  if  2  doesnot  appear  in   second  throw  and  appears  in  third  throw ∴       P o s s i b l e     o u t c o m e = 5 × 5 × 1 N o w     w e     h a v e     t h e     s e r i e s :                                             = 1 + 5 + 5 × 5 + 5 × 5 × 5 + … + 5 k − 1                                             = 1 + 5 + 5 2 + 5 3 + … + 5 k − 1                                             = 1 . ( r k − 1 ) r − 1 = 5 k − 1 5 − 1                                               = 5 k − 1 4 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y = 5 k − 1 4 .

Q:  

A die is loaded in such a way that each odd number is twice as likely to occur as each even number. Find P ( G ) , where G  is the event that a number greater than 3 occurs on a single roll of the die.

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

G i v e n     t h a t     p r o b a b i l i t y     o f     e v e n     n u m b e r s = 1 2 × p r o b a b i l i t y     o f     o d d     n u m b e r s ⇒                   P ( O d d ) : P ( E v e n ) = 2 : 1 ∴                     P ( o d d     n u m b e r ) = 2 2 + 1 = 2 3 a n d           P ( e v e n     n u m b e r ) = 1 2 + 1 = 1 3 Also,  given  that,  G  the  event  that  a  number  greater  than  3  occurs  in  a  single  throw  of  die. ∴     T h e     p o s s i b l e     o u t c o m e     a r e     4 , 5     a n d     6     o u t     o f     w h i c h     t w o     a r e     e v e n     a n d     o n e     i s     o d d . ∴     R e q u i r e d     p r o b a b i l i t y = P ( G )                                                                                                         = 2 × P ( e v e n ) × P ( o d d )                                                                                                       = 2 × 1 3 × 2 3 = 4 9 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y = 4 9 .

Q:  

In a large metropolitan area, the probabilities are 0.87, 0.36, 0.30 that a family (randomly chosen for a sample survey) owns a colour television set, a black and white television set, or both kinds of sets. What is the probability that a family owns either any one or both kinds of sets?

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

L e t     E 1     b e     t h e     e v e n t     t h a t     a     f a m i l y     o w n s     c o l o u r     t e l e v i s i o n     s e t . a n d     E 2     b e     t h e     e v e n t     t h a t     a     f a m i l y     o w n s     b l a c k     a n d     w h i t e     t e l e v i s i o n     s e t . G i v e n     t h a t     P ( E 1 ) = 0 . 8 7 ,     P ( E 2 ) = 0 . 3 6 a n d           P ( E 1 ∩ E 2 ) = 0 . 3 0 ∴ T h e     p r o b a b i l i t y     t h a t     a     f a m i l y     o w n s     e i t h e r     c o l o u r     o r     b l a c k     a n d     w h i t e     t e l e v i s i o n     s e t . ∴                   P ( E 1 ∪ E 2 ) = P ( E 1 ) + P ( E 2 ) − P ( E 1 ∩ E 2 )                                                                             = 0 . 8 7 + 0 . 3 6 − 0 . 3 0 = 0 . 9 3 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y = 0 . 9 3

Q:  

If   A and   B are mutually exclusive events, P ( A ) = 0 . 3 5  and P ( B ) = 0 . 4 5 , find:

(a) P ( A ′ )

(b) P ( B ′ )

(c) P ( A ∪ B )

(d) P ( A ∩ B )

(e) P ( A ∩ B ′ )

(f) P ( A ′ ∩ B ′ )

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

G i v e n     t h a t     P ( A ) = 0 . 3 5     a n d     P ( B ) = 0 . 4 5 Since  the  events  A  and  B  are  mutually  exclusive  then  P(A∩B)=0 ( a )                                   P ( A ' ) = 1 − P ( A ) = 1 − 0 . 3 5 = 0 . 6 5 ( b )                                   P ( B ' ) = 1 − P ( B ) = 1 − 0 . 4 5 = 0 . 5 5 ( c )                   P ( A ∪ B ) = P ( A ) + P ( B ) − P ( A ∩ B ) = 0 . 3 5 + 0 . 4 5 − 0 = 0 . 8 0 ( d )                 P ( A ∩ B ) = 0                     [ ? A     a n d     B     a r e     m u t u a l l y     e x c l u s i v e     e v e n t s ] ( e )               P ( A ∩ B ' ) = P ( A ) − P ( A ∩ B ) = 0 . 3 5 − 0 = 0 . 3 5 ( f )           P ( A ' ∩ B ' ) = 1 − P ( A ∪ B ) = 1 − 0 . 8 0 = 0 . 2 0

Q:  

A team of medical students doing their internship has to assist during surgeries at a city hospital. The probabilities of surgeries rated as very complex, complex, routine, simple, or very simple are respectively, 0.15, 0.20, 0.31, 0.26, 0.08. Find the probabilities that a particular surgery will be rated:

(a) complex or very complex;

(b) neither very complex nor very simple;

(c) routine or complex;

(d) routine or simple.

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

L e t     E 1 , E 2 , E 3 , E 4     a n d     E 5     b e     t h e     e v e n t s     t h a t     t h e     s u r g e r i e s     a r e     r a t e d     a s     v e r y     c o m p l e x ,     c o m p l e x , r o u t i n e ,     s i m p l e     a n d     v e r y     s i m p l e     r e s p e c t i v e l y . ∴ P ( E 1 ) = 0 . 1 5 ,     P ( E 2 ) = 0 . 2 0 ,     P ( E 3 ) = 0 . 3 1 ,     P ( E 4 ) = 0 . 2 6     a n d     P ( E 5 ) = 0 . 0 8 ( a ) P ( c o m p l e x     a n d     v e r y     c o m p l e x ) = P ( E 1     o r     E 2 ) ⇒ P ( E 1 ∪ E 2 ) = P ( E 1 ) + P ( E 2 ) − P ( E 1 ∩ E 2 )                                                                 = 0 . 1 5 + 0 . 2 0 − 0                                                                 = 0 . 3 5                             [ ?     A l l     e v e n t     a r e     i n d e p e n d e n t ] ( b ) P ( n e i t h e r     v e r y     c o m p l e x     n o r     v e r y     s i m p l e ) = P ( E 1 ' ∩ E 5 ' )                                                                 = P ( E 1 ∪ E 5 ) ' = 1 − P ( E 1 ∪ E 5 )                                                                 = 1 − [ P ( E 1 ) + P ( E 5 ) ] = 1 − [ 0 . 1 5 + 0 . 0 8 ]                                                                 = 1 − 0 . 2 3 = 0 . 7 7 ( c ) P ( r o u t i n e     o r     c o m p l e x ) = P ( E 3     o r     E 2 ) ⇒ P ( E 3 ∪ E 2 ) = P ( E 3 ) + P ( E 2 )                                                                 = 0 . 3 1 + 0 . 2 0 = 0 . 5 1 ( d ) P ( r o u t i n e     o r     s i m p l e ) = P ( E 3     o r     E 4 ) ⇒ P ( E 3 ∪ E 4 ) = P ( E 3 ) + P ( E 4 )                                                                 = 0 . 3 1 + 0 . 2 6 = 0 . 5 7

Q:  

Four candidates   A , B , C , D have applied for the assignment to coach a school cricket team. If   A is twice as likely to be selected as B , and   B and   C are given about the same chance of being selected, while C  is twice as likely to be selected as D , what are the probabilities that:

(a) C will be selected?

(b) A will not be selected?

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

G i v e n     t h a t     A     i s     t w i c e     a s     l i k e l y     t o     b e     s e l e c t e d     a s     B i . e . ,                                                             P ( A ) = 2 P ( B ) a n d     C     i s     t w i c e     a s     l i k e l y     t o     b e     s e l e c t e d     a s     D ∴                                                                     P ( C ) = 2 P ( D )                     ⇒ P ( B ) = 2 P ( D ) ⇒                                                               P ( A ) 2 = 2 P ( D )                     ⇒ P ( D ) = 1 4 P ( A ) N o w     B     a n d     C     a r e     g i v e n     a b o u t     t h e     s a m e     c h a n c e ∴                                                                     P ( B ) = P ( C ) Since,          sum  of  all  probabilities=1 ∴                         P ( A ) + P ( B ) + P ( C ) + P ( D ) = 1 ⇒               P ( A ) + P ( A ) 2 + P ( A ) 2 + P ( A ) 4 = 1 ⇒ 4 P ( A ) + 2 P ( A ) + 2 P ( A ) + P ( A ) = 4 ⇒                                                                                                                   9 P ( A ) = 4           ⇒ P ( A ) = 4 9 ( a )       P ( C     w i l l     b e     s e l e c t e d ) = P ( C ) = P ( B ) = P ( A ) 2                                                                                                                 = 4 9 × 1 2 = 2 9 ( b ) P ( A     w i l l     n o t     b e     s e l e c t e d ) = P ( A ' ) = 1 − P ( A )                                                                                                   &thi

Q:  

One of the four persons, John, Rita, Aslam, or Gurpreet, will be promoted next month. Consequently, the sample space consists of four elementary outcomes: 

S = {John promoted, Rita promoted, Aslam promoted, Gurpreet promoted}

You are told that the chances of John’s promotion are the same as that of Gurpreet. Rita’s chances of promotion are twice as likely as John’s. Aslam’s chances are four times that of John.

(a) Determine P (John promoted),

P (Rita promoted),

P (Aslam promoted), and

P (Gurpreet promoted).

(b) If A = {John promoted or Gurpreet promoted}, find P(A).

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

L e t     E 1 , E 2 , E 3     a n d     E 4     b e     t h e     e v e n t s     t h a t     J o h n     p r o m o t e d ,     R i t a     p r o m o t e d ,     A s l a m     p r o m o t e d a n d     G u r p r e e t     p r o m o t e d     r e s p e c t i v e l y . ∴     S a m p l e     s p a c e     S = { E 1 , E 2 , E 3 , E 4 } G i v e n     t h a t     p r o b a b i l i t y     o f     J o h n ' s     p r o m o t i o n     i s     s a m e     a s     t h a t     o f     G u r p r e e t ∴                                                   P ( E 1 ) = P ( E 4 ) R i t a ' s     c h a n c e s     o f     p r o m o t i o n     a r e     t w i c e     a s     l i k e l y     a s     J o h n ∴                                                   P ( E 2 ) = 2 P ( E 1 ) A n d     A s l a m ' s     c h a n c e s     o f     p r o m o t i o n     a r e     4     t i m e s     t h a t     o f     J o h n ∴                                                   P ( E 3 ) = 4 P ( E 1 ) Since,  the  sum  of  all  probabilities=1 ∴                 P ( E 1 ) + P ( E 2 ) + P ( E 3 ) + P ( E 4 ) = 1 ⇒ P ( E 1 ) + 2 P ( E 1 ) + 4 P ( E 1 ) + P ( E 1 ) = 1 ⇒                                                                                                                       8 P ( E 1 ) = 1 ⇒                                                                                                                             P ( E 1 ) = 1 8 ( a )     P ( J o h n     p r o m o t e d ) = P ( E 1 ) = 1 8                   P ( R i t a     p r o m o t e d ) = P ( E 2 ) = 2 P ( E 1 ) = 2 × 1 8 = 2 8           P ( A s l a m     p r o m o t e d ) = P ( E 3 ) = 4 P ( E 1 ) = 4 × 1 8 = 4 8 = 1 2 P ( G u r p r e e t     p r o m o t e d ) = P ( E 4 ) = P ( E 1 ) = 1 8 ( b ) P ( J o h n     p r o m o t e d     o r     G u r p r e e t     p r o m o t e d ) = P ( E 1 ∪ E 4 ) ⇒                       &th

Q:  

The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P (A ∩ B) = .07). Determine

(a) P (A)

(b) P (B ∩ C )

(c) P (A ∪ B)

(d) P (A ∩ B )

(e) P (B ∩ C)

(f) Probability of exactly one of the three occurs.

Read more
A: 

This is a short answer type question as classified in NCERT Exemplar

F r o m     t h e     g i v e n     V e n n     d i a g r a m ( a )                     P ( A ) = 0 . 1 3 + 0 . 0 7 = 0 . 2 0 ( b ) P ( B ∩ C ¯ ) = P ( B ) − P ( B ∩ C )                                                           = 0 . 0 7 + 0 . 1 0 + 0 . 1 5 − 0 . 1 5 = 0 . 0 7 + 0 . 1 0 = 0 . 1 7 ( c ) P ( A ∪ B ) = P ( A ) + P ( B ) − P ( A ∩ B )                                                           = 0 . 1 3 + 0 . 0 7 + 0 . 0 7 + 0 . 1 0 + 0 . 1 5 − 0 . 0 7                                                           = 0 . 1 3 + 0 . 0 7 + 0 . 1 0 + 0 . 1 5 = 0 . 4 5 ( d ) P ( A ∩ B ¯ ) = P ( A ) − P ( A ∩ B )                                                             = 0 . 1 3 + 0 . 0 7 − 0 . 0 7 = 0 . 1 3 ( e )     P ( B ∩ C ) = 0 . 1 5 ( f ) P ( e x a c t l y     o n e     o f     t h e     t h r e e     o c c u r s ) = 0 . 1 3 + 0 . 1 0 + 0 . 2 8 = 0 . 5 1

Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen Logo

Probability Long Answer Type Questions

1. One urn contains two black balls (labeled B 1 and B 1 ) and one white ball. A second urn contains one black ball and two white balls (labeled W 1 and W 2 ). Suppose the following experiment is performed: one of the two urns is chosen at random. Next, a ball is randomly chosen from the urn. Then a second ball is chosen at random from the same urn without replacing the first ball.

(a) Write the sample space showing all possible outcomes.

(b) What is the probability that two black balls are chosen?

(c) What is the probability that two balls of opposite color are chosen?

Sol. G i v e n     t h a t     o n e     o f     t h e     u r n s     i s     c h o o s e n ,     t h e n     a     b a l l     i s     r a n d o m l y     c h o o s e n     f r o m     t h e     u r n ,     t h e n a     second     b a l l     i s     c h o o s e n     a t     r a n d o m     f r o m     t h e     s a m e     u r n     w i t h o u t     r e p l a c i n g     t h e     f i r s t     b a l l ( a )     S a m p l e     s p a c e     S = { B 1 B 2 , B 1 W , B 2 B 1 , B 2 W , W B 1 , W B 2 , B W 1 , B W 2 , W 1 B , W 1 W 2 , W 2 B , W 2 W 1 }               T o t a l     n u m b e r     o f     S a m p l e     s p a c e ,     S = 1 2 ( b )     I f     t w o     b l a c k     b a l l s     a r e     c h o o s e n                 t h e n     t h e     f a v o u r a b l e     e v e n t s     a r e     B 1 B 2 , B 2 B 1     i . e .     2 ∴         R e q u i r e d     p r o b a b i l i t y = 2 1 2 = 1 6 ( c )     I f     t w o     b a l l s     o f     o p p o s i t e     c o l o u r s     a r e     c h o o s e n     t h e n ,     t h e                   r e q u i r e d     p r o b a b i l i t y = 8 1 2 = 2 3

2. A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that:

(a) All the three balls are white.

(b) All the three balls are red.

(c) One ball is red and two balls are white.

Sol. G i v e n     t h a t :                       N u m b e r     o f     r e d     b a l l s = 8                                                                 N u m b e r     o f     w h i t e     b a l l s = 5 ( a )     P ( a l l     t h e     t h r e e     b a l l s     a r e     w h i t e )                                                               = C 3 5 C 3 1 3 = 5 ! 3 ! 2 ! 1 3 ! 3 ! 1 0 ! = 5 ! 3 ! 2 ! × 3 ! 1 0 ! 1 3 !                                                                 = 5 ! 2 ! × 1 0 ! 1 3 × 1 2 × 1 1 × 1 0 !                                                                 = 5 × 4 × 3 × 2 ! 2 ! × 1 1 3 × 1 2 × 1 1 = 5 × 4 × 3 1 3 × 1 2 × 1 1 = 5 1 4 3 ( b )     P ( a l l     t h e     t h r e e     b a l l s     a r e     r e d )                                                               = C 3 8 C 3 1 3 = 8 ! 3 ! 5 ! 1 3 ! 3 ! 1 0 ! = 8 ! 3 ! 5 ! × 3 ! 1 0 ! 1 3 !                                                                 = 8 × 7 × 6 × 5 ! 5 ! × 1 0 ! 1 3 × 1 2 × 1 1 × 1 0 !                                                                 = 8 × 7 × 6 1 3 × 1 2 × 1 1 = 2 8 1 4 3 ( c )     P ( o n e     b a l l     i s     r e d     a n d     t w o     b a l l s     a r e     w h i t e )                                                               = C 1 8 × C 2 5 C 3 1 3 = 8 ! 7 ! 1 ! × 5 ! 2 ! 3 ! 1 3 ! 3 ! 1 0 ! = 8 × 7 ! 7 ! × 5 × 4 × 3 × 2 ! 2 ! × 3 ! × 3 !   × 1 0 ! 1 3 × 1 2 × 1 1 × 1 0 !                                                               = 8 × 5 1 3 × 1 1 = 4 0 1 4 3
Q&A Icon
Commonly asked questions
Q:  

One urn contains two black balls (labeled B 1 and B 1 ) and one white ball. A second urn contains one black ball and two white balls (labeled W 1 and W 2 ). Suppose the following experiment is performed: one of the two urns is chosen at random. Next, a ball is randomly chosen from the urn. Then a second ball is chosen at random from the same urn without replacing the first ball.

(a) Write the sample space showing all possible outcomes.

(b) What is the probability that two black balls are chosen?

(c) What is the probability that two balls of opposite color are chosen?

Read more
A: 

This is a long answer type question as classified in NCERT Exemplar

G i v e n     t h a t     o n e     o f     t h e     u r n s     i s     c h o o s e n ,     t h e n     a     b a l l     i s     r a n d o m l y     c h o o s e n     f r o m     t h e     u r n ,     t h e n a  second  ball  is  choosen  at  random  from  the  same  urn  without  replacing  the  first  ball ( a )     S a m p l e     s p a c e     S = { B 1 B 2 , B 1 W , B 2 B 1 , B 2 W , W B 1 , W B 2 , B W 1 , B W 2 , W 1 B , W 1 W 2 , W 2 B , W 2 W 1 }               T o t a l     n u m b e r     o f     S a m p l e     s p a c e ,     S = 1 2 ( b )     I f     t w o     b l a c k     b a l l s     a r e     c h o o s e n                 t h e n     t h e     f a v o u r a b l e     e v e n t s     a r e     B 1 B 2 , B 2 B 1     i . e .     2 ∴         R e q u i r e d     p r o b a b i l i t y = 2 1 2 = 1 6 ( c )     I f     t w o     b a l l s     o f     o p p o s i t e     c o l o u r s     a r e     c h o o s e n     t h e n ,     t h e                   r e q u i r e d     p r o b a b i l i t y = 8 1 2 = 2 3

Q:  

A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that:

(a) All the three balls are white.

(b) All the three balls are red.

(c) One ball is red and two balls are white.

Read more
A: 

This is a long answer type question as classified in NCERT Exemplar

G i v e n     t h a t :                       N u m b e r     o f     r e d     b a l l s = 8                                                                 N u m b e r     o f     w h i t e     b a l l s = 5 ( a )     P ( a l l     t h e     t h r e e     b a l l s     a r e     w h i t e )                                                               = C 3 5 C 3 1 3 = 5 ! 3 ! 2 ! 1 3 ! 3 ! 1 0 ! = 5 ! 3 ! 2 ! × 3 ! 1 0 ! 1 3 !                                                                 = 5 ! 2 ! × 1 0 ! 1 3 × 1 2 × 1 1 × 1 0 !                                                                 = 5 × 4 × 3 × 2 ! 2 ! × 1 1 3 × 1 2 × 1 1 = 5 × 4 × 3 1 3 × 1 2 × 1 1 = 5 1 4 3 ( b )     P ( a l l     t h e     t h r e e     b a l l s     a r e     r e d )                                                               = C 3 8 C 3 1 3 = 8 ! 3 ! 5 ! 1 3 ! 3 ! 1 0 ! = 8 ! 3 ! 5 ! × 3 ! 1 0 ! 1 3 !                                                                 = 8 × 7 × 6 × 5 ! 5 ! × 1 0 ! 1 3 × 1 2 × 1 1 × 1 0 !                                                                 = 8 × 7 × 6 1 3 × 1 2 × 1 1 = 2 8 1 4 3 ( c )     P ( o n e     b a l l     i s     r e d     a n d     t w o     b a l l s     a r e     w h i t e )                                                               = C 1 8 × C 2 5 C 3 1 3 = 8 ! 7 ! 1 ! × 5 ! 2 ! 3 ! 1 3 ! 3 ! 1 0 ! = 8 × 7 ! 7 ! × 5 × 4 × 3 × 2 ! 2 ! × 3 ! × 3 !   × 1 0 ! 1 3 × 1 2 × 1 1 × 1 0 !                                                               = 8 × 5 1 3 × 1 1 = 4 0 1 4 3

Q:  

If the letters of the word ASSASSINATION are arranged at random, find the probability that:

(a) Four S’s come consecutively in the word.

(b) Two I’s and two N’s come together.

(c) All A’s are not coming together.

(d) No two A’s are coming together.

Read more
A: 

This is a long answer type question as classified in NCERT Exemplar

T o t a l     n u m b e r     o f     w o r d     i s     A S S A S S I N A T I O N     a r e     1 3 . W h e r e ,     w e     h a v e     3 A ' s ,     4 S ' s ,     2 I ' s ,     2 N ' s ,     1 T ' s     a n d     1 O ' s . (a)  If  4S's  come  consecutively  in  the  word,  then  arrangement  may  be  as  follows:                                   S S S S 1     G r o u p     A A A I I N N T O ? 9     o t h e r s ∴           N u m b e r     o f     w o r d s     w h e n     a l l     S ' s     a r e     t o g e t h e r = 1 0 ! 3 ! 2 ! 2 !

              a n d     t h e     t o t a l     n u m b e r     o f     w o r d     f o r m e d     f r o m     t h e     w o r d s     A S S A S S I N A T I O N = 1 3 ! 3 ! 4 ! 2 ! 2 ! ∴           R e q u i r e d     p r o b a b i l i t y = 1 0 ! 3 ! 2 ! 2 ! 1 3 ! 3 ! 4 ! 2 ! 2 ! = 1 0 ! 3 ! 2 ! 2 ! × 3 ! 4 ! 2 ! 2 ! 1 3 !                                                                                                                 = 1 0 ! 4 ! 1 3 ! = 1 0 !   × 4 × 3 × 2 1 3 × 1 2 × 1 1 × 1 0 ! = 2 1 4 3 ( b )     I f     2 I ' s     a n d     2 N ' s     c o m e     t o g e t h e r     t h e n     t h e r e     a r e     1 0     a l p h a b e t s                 N u m b e r     o f     w o r d s     w h e n     2 I ' s     a n d     2 N ' s     c o m e     t o g e t h e r = 1 0 ! 3 ! 4 ! × 4 ! 2 ! 2 ! ∴           R e q u i r e d     p r o b a b i l i t y = 1 0 ! 3 ! 4 ! × 4 ! 2 ! 2 ! 1 3 ! 3 ! 4 ! 2 ! 2 ! = 1 0 ! 3 ! 4 ! × 4 ! 2 ! 2 ! × 3 ! 4 ! 2 ! 2 ! 1 3 !                                                                                                                 = 4 ! 1 0 ! 1 3 ! = 4 × 3 × 2 × 1 0 ! 1 3 × 1 2 × 1 1 × 1 0 ! = 2 1 4 3 (c)  If  all  A's  are  coming  together,  then  there  are  11  alphabets               N u m b e r     o f     w o r d s     w h e n     a l l     A ' s     c o m e     t o g e t h e r = 1 1 ! 4 ! 2 ! 2 ! ∴           P r o b a b i l i t y     w h e n     a l l     A ' s     c o m e     t o g e t h e r = 1 1 ! 4 ! 2 ! 2 ! 1 3 ! 4 ! 3 ! 2 ! 2 ! = 1 1 ! 4 ! 2 ! 2 ! × 4 ! 3 ! 2 ! 2 ! 1 3 !                                                                                                                                                                                                   = 1 1 ! 3 ! 1 3 ! = 6 1 3 × 1 2 = 1 2 6 ∴       R e q u i r e d     p r o b a b i l i t y     w h e n     a l l     A ' s     d o     n o t     c o m e     t o g e t h e r = 1 − 1 2 6 = 2 5 2 6 ( d )     I f     n o     t w o     A ' s     a r e     t o g e t h e r ,     t h e n     a r r a n g i n g     t h e     a l p h a b e t s     e x c e p t     A ' s                   − S − S − S − S − I − N − T − I − O − N −                   N u m b e r     o f     w a y s     o f     a r r a n g i n g     a l l     a l p h a b e t s       e x c e p t     A ' s = 1 0 ! 4 ! 2 ! 2 !                   T h e r e     a r e     1 1     v a c a n t     p l a c e s     b e t w e e n     t h e s e     a l p h a b e t s . ∴               3 A ' s     c a n     b e     p l a c e d     i n     1 1     p l a c e s     i n     C 3 1 1     w a y s = 1 1 ! 3 ! 8 ! ∴             T o t a l     n u m b e r     o f     w o r d s     w h e n     n o     t w o     A ' s     c o m e     t o g e t h e r = 1 1 ! 3 ! 8 ! × 1 0 ! 4 ! 2 ! 2 ! ∴           R e q u i r e d     p r o b a b i l i t y = 1 1 ! 3 ! 8 ! × 1 0 ! 4 ! 2 ! 2 ! × 4 ! 3 ! 2 ! 2 ! 1 3 !                                                                                                               = 1 0 ! 8 !   × 1 3 × 1 2 = 1 0 × 9 × 8 ! 8 !   × 1 3 × 1 2 = 1 0 × 9 1 3 × 1 2 = 1 5 2 6

Q:  

A card is drawn from a deck of 52 cards. Find the probability of getting a king or a heart or a red card.

Read more
A: 

This is a long answer type question as classified in NCERT Exemplar

T o t a l     n u m b e r     o f     c a r d s = 5 2 F a v o u r a b l e     e v e n t s = 4     k i n g s + 1 3     h e a r t s + 2 6     r e d − 1 3 − 2 = 2 8 ∴     R e q u i r e d     p r o b a b i l i t y = 2 8 5 2 = 7 1 3 .

Q:  

A sample space consists of 9 elementary outcomes   e 1 , e 2 , … , e 9 whose probabilities are:

  P ( e 1 ) = P ( e 2 ) = 0 . 0 8 , P ( e 3 ) = P ( e 4 ) = P ( e 5 ) = 0 . 1 , P ( e 6 ) = P ( e 7 ) = 0 . 2 , P ( e 8 ) = P ( e 9 ) = 0 . 0 7 .

Suppose  A = { e 1 , e 5 , e 8 } , B = { e 2 , e 5 , e 8 , e 9 } .

(a) Calculate P ( A ) , P ( B ) , and P ( A ∩ B ) .

(b) Using the addition law of probability, calculate P ( A ∪ B ) .

(c) List the composition of the event A ∪ B , and calculate   P ( A ∪ B ) by adding the probabilities of the elementary outcomes.

(d) Calculate   P ( B ′ ) from P ( B ) , also calculate   P ( B ′ ) directly from the elementary outcomes of B .

Read more
A: 

This is a long answer type question as classified in NCERT Exemplar

G i v e n     t h a t :             S = { e 1 , e 2 , e 3 , e 4 , e 5 , e 6 , e 7 , e 8 , e 9 }                                                             A = { e 1 , e 5 , e 5 }     a n d     B = { e 2 , e 5 , e 5 , e 9 }                                             P ( e 1 ) = P ( e 2 ) = 0 . 0 8                                             P ( e 3 ) = P ( e 4 ) = P ( e 5 ) = 0 . 1                                             P ( e 6 ) = P ( e 7 ) = 0 . 2                                             P ( e 8 ) = P ( e 9 ) = 0 . 0 7 ( a )                                 P ( A ) = P ( e 1 ) + P ( e 5 ) + P ( e 8 )                                                                       = 0 . 0 8 + 0 . 1 + 0 . 0 7 = 0 . 2 5 ( b )               P ( A ∪ B ) = P ( A ) + P ( B ) − P ( A ∩ B )                                                                 … ( i )               B u t                 P ( B ) = P ( e 2 ) + P ( e 5 ) + P ( e 8 ) + P ( e 9 )                                                                         = 0 . 0 8 + 0 . 1 + 0 . 0 7 + 0 . 0 7 = 0 . 3 2       a n d     P ( A ∩ B ) = { e 5 , e 8 } = P ( e 5 ) + P ( e 8 )                                                                         = 0 . 1 + 0 . 0 7 = 0 . 1 7         P u t t i n g     t h e     v a l u e s     i n     e q n . ( i )     w e     g e t         P ( A ∪ B ) = 0 . 2 5 + 0 . 3 2 − 0 . 1 7 = 0 . 4 0 ( c )         P ( A ∪ B ) = { e 1 , e 2 , e 5 , e 8 , e 9 } ∴               P ( A ∪ B ) = P ( e 1 ) + P ( e 2 ) + P ( e 5 ) + P ( e 8 ) + P ( e 9 )                                                                 = 0 . 0 8 + 0 . 0 8 + 0 . 1 + 0 . 0 7 + 0 . 0 7 = 0 . 4 0 ( d )     &t

Q:  

Determine the probability p, for each of the following events.

(a) An odd number appears in a single toss of a fair die.

(b) At least one head appears in two tosses of a fair coin.

(c) A king, 9 of hearts, or 3 of spades appears in drawing a single card from a well shuffled ordinary deck of 52 cards.

(d) The sum of 6 appears in a single toss of a pair of fair dice.

Read more
A: 

This is a long answer type question as classified in NCERT Exemplar

(a)  Possible  outcome  of  a  single  throw  of  die                 S = { 1 , 2 , 3 , 4 , 5 , 6 }     o u t     o f     w h i c h     1 , 3 , 5     a r e     o d d ∴           R e q u i r e d     p r o b a b i l i t y = 3 6 = 1 2 ( b )     W h e n     a     f a i r     c o i n     i s     t o s s e d     t w i c e ,     t h e n     t h e     s a m p l e     s p a c e                 S = { H H , H T , T H , T T } ∴           P r o b a b i l i t y     o f     g e t t i n g     a t l e a s t     o n e     h e a d ( H H , H T , T H ) = 3 4 ( c )     F a v o u r a b l e     e v e n t s     a r e     4     k i n g s + 2     o f     h e a r t s + 3     o f     s p a d e s = 4 + 1 + 1 = 6 = 6 5 2 = 3 2 6 ( d )     W h e n     a     p a i r     o f     d i c e     i s     r o l l e d ,     t h e n     t o t a l     n u m b e r     o f     s a m p l e     s p a c e = 3 6     o u t     o f     w h i c h                 ( 1 , 5 ) , ( 5 , 1 ) , ( 2 , 4 ) , ( 4 , 2 )     a n d     ( 3 , 3 )     a r e     t h e     f a v o u r a b l e     e v e n t s ∴             R e q u i r e d     p r o b a b i l i t y = 5 3 6 .

Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen Logo

Probability Objective Type Questions

1. In a non-leap year, the probability of having 53 Tuesdays or 53 Wednesdays is:

(a) 1 7

(b) 2 7

(c) 3 7

(d) None of these

Sol. T h e r e     a r e     3 6 5     d a y s     i n     a     l e a p     y e a r     a n d     t h e r e     a r e     7     d a y s     i n     a     w e e k . ∴     3 6 5 ÷ 7 = 5 2     w e e k s + 1     d a y S o ,     t h i s     d a y     m a y     b e     T u e s d a y     o r     W e d n e s d a y . S o ,     t h e     r e q u i r e d     p r o b a b i l i t y = 1 7 . H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( a ) .

2. Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive:

(a) 1 8 6 1 9 0  

(b) 1 8 7 1 9 0  

(c) 1 8 8 1 9 0  

(d) 18/20C3

Sol. S e t     o f     t h r e e     c o n s e c u t i v e     n u m b e r s     f r o m     1     t o     2 0     a r e     1 , 2 , 3 ;     2 , 3 , 4 ;     3 , 4 , 5 ;     … ; 1 8 , 1 9 , 2 0 . S o ,     t h e     p r o b a b i l i t y     t h a t     t h e     n u m b e r s     a r e     consecutive                                                 = 1 8 C 3 2 0 = 1 8 2 0 ! 3 ! 7 ! = 1 8 . 3 ! 7 ! 2 0 !                                                 = 1 8 × 3 × 2 × 1 7 ! 2 0 × 1 9 × 1 8 × 1 7 ! = 3 × 2 2 0 × 1 9 = 3 1 9 0 ∴     P ( t h r e e     n u m b e r s     a r e     n o t     consecutive ) = 1 − 3 1 9 0 = 1 8 7 1 9 0 H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( b ) .
Q&A Icon
Commonly asked questions
Q:  

A single letter is selected at random from the word “PROBABILITY.” The probability that it is a vowel is:

(a) 1 3

(b)  14

(c) 2 1 1  

(d) None of these.

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

Total  number  of  alphabets  in  probability=11Number  of  vowels=4∴   Required probability=411Hence,   the  correct  option  is   (b).

Q:  

In a non-leap year, the probability of having 53 Tuesdays or 53 Wednesdays is:

(a) 1 7

(b) 2 7

(c) 3 7

(d) None of these

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

T h e r e     a r e     3 6 5     d a y s     i n     a     l e a p     y e a r     a n d     t h e r e     a r e     7     d a y s     i n     a     w e e k . ∴     3 6 5 ÷ 7 = 5 2     w e e k s + 1     d a y S o ,     t h i s     d a y     m a y     b e     T u e s d a y     o r     W e d n e s d a y . S o ,     t h e     r e q u i r e d     p r o b a b i l i t y = 1 7 . H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( a ) .

Q:  

Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive:

(a) 1 8 6 1 9 0  

(b) 1 8 7 1 9 0  

(c) 1 8 8 1 9 0  

(d) 18/20C3

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

S e t     o f     t h r e e     c o n s e c u t i v e     n u m b e r s     f r o m     1     t o     2 0     a r e     1 , 2 , 3 ;     2 , 3 , 4 ;     3 , 4 , 5 ;     … ; 1 8 , 1 9 , 2 0 . So,  the  probability  that  the  numbers  are  consecutive                                                 = 1 8 C 3 2 0 = 1 8 2 0 ! 3 ! 7 ! = 1 8 . 3 ! 7 ! 2 0 !                                                 = 1 8 × 3 × 2 × 1 7 ! 2 0 × 1 9 × 1 8 × 1 7 ! = 3 × 2 2 0 × 1 9 = 3 1 9 0 ∴     P ( three  numbers  are  not  consecutive ) = 1 − 3 1 9 0 = 1 8 7 1 9 0 H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( b ) .

Q:  

While shuffling a pack of 52 playing cards, 2 are accidentally dropped. Find the probability that the missing cards are of different colors

(a) 2 9 5 2    

(b) 1 2  

(c) 2 6 5 1  

(d) 2 7 5 1  

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

W e     k n o w     t h a t     o u t     o f     5 2     p l a y i n g     c a r d s     2 6     a r e     o f     r e d     a n d     2 6     a r e     o f     b l a c k     c o l o u r . ∴     P ( b o t h     c a r d s     o f     d i f f e r e n t     c o l o u r ) = 2 6 5 2 × 2 6 5 1 + 2 6 5 2 × 2 6 5 1                                                                                                                                                               = 2 × 2 6 5 2 × 2 6 5 1 = 2 6 5 1 H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( c ) .

Q:  

Seven persons are to be seated in a row. The probability that two particular persons sit next to each other is:

(a) 1 3  

(b) 1 6  

(c)   2 7  

(d) 1 2  

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

The  two  particular  persons  to  be  seated  next  each  other  then,  they  form  one  group.Now  the  permutation  of  6  persons=6!  ×  2!and  Total  number  of  permutation  of  7  persons=7!∴   Required probability=6!  ×  2!7!                                                    =6!  ×  27×6!=27Hence,  the  correct  option  is  (c).

Q:  

If P(A ∪ B) = P(A ∩ B) for any two events A and B, then:

(a) P(A) = P(B)

(b) P(A) > P(B)

(c) P(A) < P(B)

(d) None of these.

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

G i v e n     t h a t :                 P ( A ∪ B ) = P ( A ∩ B ) ⇒                                                                         P ( A ) + P ( B ) − P ( A ∩ B ) = P ( A ∩ B ) ⇒ [ P ( A ) − P ( A ∩ B ) ] + [ P ( B ) − P ( A ∩ B ) ] = 0 B u t     P ( A ) − P ( A ∩ B ) ≥ 0                   [ ? P ( A ∩ B ) ≤ P ( A )     o r     P ( B ) ]                                                     … ( i ) a n d     P ( B ) − P ( A ∩ B ) ≥ 0                                                                                                                                                                                                           … ( i i ) F r o m     e q n . ( i )     a n d     ( i i )     w e     g e t                           P ( A ) = P ( B ) H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( a ) .

Q:  

Six boys and six girls sit in a row at random. The probability that all the girls sit together is:

(a) 1 4 3 2

(b) 1 2 4 3 1

(c) 1 1 3 2

(d) None of these.

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

If  all  the  girls  sit  together,  then  we  consider  it  as  1  group∴  Total  number  of  arrangement  of  6+1=7  persons  in  a  row=7!and  the  girls  also    their  places  with  6!  ways.∴    Required probability=6! 7!12!=6×5×4×3×2×7!12×11×10×9×8×7!=1132Hence,  the  correct  option  is  (c).

.

Q:  

If the probabilities for A to fail in an examination are 0.2 and for B are 0.3, then the probability that either A or B fails is:

(a) > 0.5

(b) 0.5

(c) ≤ 0.5

(d) 0

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

Given  that:           P(A  fails)=0.2                                    P(B  fails)=0.3∴        P(either  A  or  B  fails)≤P(A  fails)+P(B  fails)                                                              ≤0.2+0.3                                                              ≤0.5Hence,  the  correct  option  is  (c).

Q:  

The probability that at least one of the events A and B occurs is 0.6. If A and B occur simultaneously with a probability of 0.2, then P(A) + P(B) is:

(a) 0.4

(b) 0.8

(c) 1.2

(d) 1.6

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

Given  that:           P(A∪B)=0.6                                    P(A∩B)=0.2∴          P(A∪B)=P(A)+P(B)−P(A∩B)⇒                       0.6=P(A)+P(B)−0.2⇒P(A)+P(B)=0.6+0.2=0.8and        1−P(A¯)+1−P(B¯)=0.8⇒     P(A¯)+P(B¯)=2−0.8=1.2Hence,  the  correct  option  is  (c).

Q:  

If M and N are any two events, the probability that at least one of them occurs is:

(a) P(M) + P(N) – 2P(M ∩ N)

(b) P(M) + P(N) – P(M ∩ N)

(c) P(M) + P(N) + P(M ∩ N)

(d) P(M) + P(N) + 2P(M ∩ N)

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

If  M  and  N  are  any  two  events,   then          P (M∪N)=P (M)+P (N)−P (M∩N)Hence,   the  correct  option  is   (b).

Q:  

Without repetition of the numbers, four-digit numbers are formed with the numbers 0, 2, 3, 5. The probability of such a number divisible by 5 is:

(a) 1 5  

(b) 4 5

(c) 1 3 0  

(d) 5 9

Read more
A: 

This is an Objective Type Questions as classified in NCERT ExemplarFour  digit  number    the  digits  0,2,3,5  with  out  repetition  and  divisible  by  5  with  thegiven  condition  isIf  unit  place  be  filled  with  0Then  the  number  of  ways=3×2×1×1=6If  unit  place  be  filled  with  5Then  the  number  of  ways=2×2×1×1=4∴  Total  number  of  ways=6+4=10Total  number  of  ways  of  arranging  the  digits  0,2,3,5  to  form  4digit  numbers  withoutrepetition  is  3×3×2×1=18∴   Required probability=1018=59Hence,  the  correct  option  is  (d).

 



Q:  

If A and B are mutually exclusive events, then: (A) P

(a) ≤ P(B)

(b) P(A) ≥ P(B)

(c) P(A) < P(B)

(d) None of these.

Read more
A: 

This is an Objective Type Questions as classified in NCERT Exemplar

F o r     m u t u a l l y     e x c l u s i v e     e v e n t s ,                     P ( A ∩ B ) = 0 ∴               P ( A ∪ B ) = P ( A ) + P ( B )                 [ ? P ( A ∩ B ) = 0 ] ⇒                 P ( A ) + P ( B ) ≤ 1 ⇒ P ( A ) + 1 − P ( B ¯ ) ≤ 1                                                 [ ? P ( B ) = 1 − P ( B ¯ ) ] ⇒                 P ( A ) − P ( B ¯ ) ≤ 0 ⇒                                                 P ( A ) ≤ P ( B ¯ ) H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( a ) .

Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen Logo

Probability True or False Type Questions

1. The probability that a person visiting a zoo will see the giraffe is 0.72, the probability that he will see the bears is 0.84, and the probability that he will see both is 0.52.

Sol. G i v e n     t h a t :                                                                   P ( t o     s e e     g i r a f f e e ) = 0 . 7 2                                                                               P ( t o     s e e     b e a r ) = 0 . 8 4 P ( t o     s e e     b o t h     g i r a f f e e     a n d     b e a r ) = 0 . 5 2 ∴                         P ( t o     s e e     g i r a f f e e     o r     b e a r ) = P ( t o     s e e     g i r a f f e e ) + P ( t o     s e e     b e a r ) − P ( t o     s e e     b o t h )                                                                                                                                                   = 0 . 7 2 + 0 . 8 4 − 0 . 5 2                                                                                                                                                   = 1 . 0 4     w h i c h     i s     n o t     p o s s i b l e . H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .

2. The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get a compartment is 0.96.

Sol. L e t     E     b e     t h e     e v e n t     t h a t     t h e     s t u d e n t     w i l l     p a s s a n d     F     b e     t h e     e v e n t     t h a t     h e     w i l l     g e t     c o m p a r t m e n t ∴ P ( E ) = 0 . 7 3 , P ( F ) = 0 . 1 3     a n d     P ( E ∪ F ) = 0 . 9 6 ∴                   P ( E ∪ F ) = P ( E ) + P ( F ) − P ( E ∩ F )                                                                     = 0 . 7 3 + 0 . 1 3 − 0                       [ ∵ P ( E ∩ F ) = 0 ]                                                                     = 0 . 8 6 B u t         P ( E ∪ F ) = 0 . 9 6 H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .
Q&A Icon
Commonly asked questions
Q:  

The probability that a person visiting a zoo will see the giraffe is 0.72, the probability that he will see the bears is 0.84, and the probability that he will see both is 0.52.

Read more
A: 

This is a True or False Type Questions as classified in NCERT Exemplar

G i v e n     t h a t :                                                                   P ( t o     s e e     g i r a f f e e ) = 0 . 7 2                                                                               P ( t o     s e e     b e a r ) = 0 . 8 4 P ( t o     s e e     b o t h     g i r a f f e e     a n d     b e a r ) = 0 . 5 2 ∴                         P ( t o     s e e     g i r a f f e e     o r     b e a r ) = P ( t o     s e e     g i r a f f e e ) + P ( t o     s e e     b e a r ) − P ( t o     s e e     b o t h )                                                                                                                                                   = 0 . 7 2 + 0 . 8 4 − 0 . 5 2                                                                                                                                                   = 1 . 0 4     w h i c h     i s     n o t     p o s s i b l e . H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .

Q:  

The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get a compartment is 0.96.

Read more
A: 

This is a True or False Type Questions as classified in NCERT Exemplar

L e t     E     b e     t h e     e v e n t     t h a t     t h e     s t u d e n t     w i l l     p a s s a n d     F     b e     t h e     e v e n t     t h a t     h e     w i l l     g e t     c o m p a r t m e n t ∴ P ( E ) = 0 . 7 3 , P ( F ) = 0 . 1 3     a n d     P ( E ∪ F ) = 0 . 9 6 ∴                   P ( E ∪ F ) = P ( E ) + P ( F ) − P ( E ∩ F )                                                                     = 0 . 7 3 + 0 . 1 3 − 0                       [ ? P ( E ∩ F ) = 0 ]                                                                     = 0 . 8 6 B u t         P ( E ∪ F ) = 0 . 9 6 H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .

Q:  

The probabilities that a typist will make 0, 1, 2, 3, 4, or 5 or more mistakes in typing a report are, respectively, 0.12, 0.25, 0.36, 0.14, 0.08, 0.11.

Read more
A: 

This is a True or False Type Questions as classified in NCERT Exemplar

S u m     o f     a l l     p r o b a b i l i t i e s = 1 ∴ P ( 0 ) + P ( 1 ) + P ( 2 ) + P ( 3 ) + P ( 4 ) + P ( 5 ) = 0 . 1 2 + 0 . 2 5 + 0 . 3 6 + 0 . 1 4 + 0 . 0 8 + 0 . 1 1                                                                                                                                                                                       = 1 . 0 6 > 1 H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .

Q:  

If A and B are two candidates seeking admission in an engineering college, the probability that A is selected is 0.5, and the probability that both A and B are selected is at most 0.3. Is it possible that the probability of B getting selected is 0.7?

Read more
A: 

This is a True or False Type Questions as classified in NCERT Exemplar

G i v e n     t h a t     P ( A ) = 0 . 5     a n d     P ( A ∩ B ) ≤ 0 . 3 N o w     P ( A ) × P ( B ) ≤ 0 . 3 ⇒                           0 . 5 × P ( B ) ≤ 0 . 3 ⇒                                               P ( B ) ≤ 0 . 3 0 . 5 ⇒                                               P ( B ) ≤ 0 . 6 H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e '

Q:  

The probability of the occurrence of event A is 0.7 and that of the occurrence of event B is 0.3, but the probability of the occurrence of both is 0.4.

Read more
A: 

This is a True or False Type Questions as classified in NCERT Exemplar

H e r e ,     P ( A ) = 0 . 7 ,     P ( B ) = 0 . 3 ∴     P ( A ∩ B ) = P ( A ) × P ( B )                                                       = 0 . 7 × 0 . 3 = 0 . 2 1 B u t     t h e     g i v e n     p r o b a b i l i t y     i s     0 . 4 H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e '

Q:  

The probability of the intersection of two events A and B is always less than or equal to the probability of A.

Read more
A: 

This is a True or False Type Questions as classified in NCERT Exemplar

H e r e ,     P ( A ∩ B ) ≤ P ( A )     W h i c h     i s     a l w a y s     t r u e . H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' T r u e '

Q:  

The sum of probabilities of two students getting distinction in their final examinations is 1.2.

A: 

This is a True or False Type Questions as classified in NCERT Exemplar

Since,  the  two  given  events  are  not  related  to  the  same  Sample  space.∴The  sum  of  probabilities  of  two  students  getting  distinction  in  their  final      may  be  1.2Hence,  the  given  statement  is  'True'

Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen Logo

Probability Fill in the blank Type Questions

1. The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is ___.
Sol. P ( loosing     t h e     g a m e ) = 1 − ( 0 . 7 7 + 0 . 0 8 )                                                                                             = 1 − 0 . 8 5 = 0 . 1 5 H e n c e ,     t h e     v a l u e     o f     t h e     f i l l e r     i s     0 . 1 5

2. If e₁, e₂, e₃, e₄ are the four elementary outcomes in a sample space and P(e₁) = 0.1, P(e₂) = 0.5, P(e₃) = 0.1, then the probability of e₄ is __.

Sol. W e     k n o w     t h a t     t h e     s u m     o f     a l l     p r o b a b i l i t i e s = 1 ∴                         P ( e 1 ) + P ( e 2 ) + P ( e 3 ) + P ( e 4 ) = 1 ⇒                                                         0 . 1 + 0 . 5 + 0 . 1 + P ( e 4 ) = 1 ⇒                                                                                                     0 . 7 + P ( e 4 ) = 1 ⇒                       P ( e 4 ) = 1 − 0 . 7 = 0 . 3 H e n c e ,     t h e     v a l u e     o f     t h e     f i l l e r     i s     0 . 3
Q&A Icon
Commonly asked questions
Q:  

The probability that a person visiting a zoo will see the giraffe is 0.72, the probability that he will see the bears is 0.84, and the probability that he will see both is 0.52.

Read more
A: 

This is a Fill in the blanks Type Questions as classified in NCERT Exemplar

G i v e n     t h a t :                                                                   P ( t o     s e e     g i r a f f e e ) = 0 . 7 2                                                                               P ( t o     s e e     b e a r ) = 0 . 8 4 P ( t o     s e e     b o t h     g i r a f f e e     a n d     b e a r ) = 0 . 5 2 ∴                         P ( t o     s e e     g i r a f f e e     o r     b e a r ) = P ( t o     s e e     g i r a f f e e ) + P ( t o     s e e     b e a r ) − P ( t o     s e e     b o t h )                                                                                                                                                   = 0 . 7 2 + 0 . 8 4 − 0 . 5 2                                                                                                                                                   = 1 . 0 4     w h i c h     i s     n o t     p o s s i b l e . H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .

Q:  

The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get a compartment is 0.96.

Read more
A: 

This is a Fill in the blanks Type Questions as classified in NCERT Exemplar

L e t     E     b e     t h e     e v e n t     t h a t     t h e     s t u d e n t     w i l l     p a s s a n d     F     b e     t h e     e v e n t     t h a t     h e     w i l l     g e t     c o m p a r t m e n t ∴ P ( E ) = 0 . 7 3 , P ( F ) = 0 . 1 3     a n d     P ( E ∪ F ) = 0 . 9 6 ∴                   P ( E ∪ F ) = P ( E ) + P ( F ) − P ( E ∩ F )                                                                     = 0 . 7 3 + 0 . 1 3 − 0                       [ ? P ( E ∩ F ) = 0 ]                                                                     = 0 . 8 6 B u t         P ( E ∪ F ) = 0 . 9 6 H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .

Q:  

If A and B are two events associated with a random experiment such that P(A) = 0.3, P(B) = 0.2, and P(A ∩ B) = 0.1, then the value of P(A ∩ B') is ___.

Read more
A: 

This is a Fill in the blanks Type Questions as classified in NCERT Exemplar

G i v e n     t h a t :                         P ( A ) = 0 . 3 ,     P ( B ) = 0 . 2 a n d                                       P ( A ∩ B ) = 0 . 1 ∴                                                   P ( A ∩ B ¯ ) = P ( A ) − P ( A ∩ B )                                                                                                     = 0 . 3 − 0 . 1 = 0 . 2 H e n c e ,     t h e     v a l u e     o f     t h e     f i l l e r     i s     0 . 2

Q:  

The probability of happening of event A is 0.5, and that of B is 0.3. If A and B are mutually exclusive events, then the probability of neither A nor B is ____.

Read more
A: 

This is a Fill in the blanks Type Questions as classified in NCERT Exemplar

G i v e n     t h a t :                         P ( A ) = 0 . 5 ,     P ( B ) = 0 . 3 a n d                                       P ( A ∩ B ) = 0                 [ ? A     a n d     B     a r e     m u t u a l l y     e x c l u s i v e     e v e n t s ] ∴                                                   P ( A ¯ ∩ B ¯ ) = P ( A ∪ B ¯ )                                                                                                     = 1 − P ( A ∪ B ) = 1 − [ P ( A ) + P ( B ) ]                                                                                                     = 1 − ( 0 . 5 + 0 . 3 ) = 1 − 0 . 8 = 0 . 2 H e n c e ,     t h e     v a l u e     o f     t h e     f i l l e r     i s     0 . 2

Q:  

Let S = {1, 2, 3, 4, 5, 6} and E = {1, 3, 5}. Then E is ___.

A: 

This is a Fill in the blanks Type Questions as classified in NCERT Exemplar

G i v e n     t h a t :                           S = { 1 , 2 , 3 , 4 , 5 , 6 }                                                                           E = { 1 , 3 , 5 } ∴                                                                       E ¯ = S − E = { 2 , 4 , 6 } H e n c e ,     t h e     v a l u e     o f     t h e     f i l l e r     i s     { 2 , 4 , 6 } .

Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen Logo

Probability Matching Type Questions

1. Match the proposed probability under Column C1 with the appropriate written description under column C2 :

C1                                                                    C2

Probability                                                       Written Description

(a) 0.95                                                             (i) An incorrect assignment

(b) 0.02                                                             (ii) No chance of happening

(c) – 0.3                                                            (iii) As much chance of happening as not.

(d) 0.5                                                               (iv) Very likely to happen

(e) 0                                                                  (v) Very little chance of happening

Sol. ( a ) 0 . 9 5 = V e r y     l i k e l y     t o     h a p p e n ,     s o     i t     i s     c l o s e     t o     1 . ( b ) 0 . 0 2 = V e r y     l i t t l e     c h a n c e     o f     h a p p e n i n g     a s     t h e     p r o b a b i l i t y     i s     v e r y     l o w . ( c ) − 0 . 3 = A n     i n c o r r e c t     a s s i g n m e n t     b e c a u s e     p r o b a b i l i t y     i s     n e v e r     n e g a t i v e . ( d )         0 . 5 = A s     m u c h     c h a n c e     o f     h a p p e n i n g     a s     n o t     b e c a u s e     s u m     o f     c h a n c e s     o f     h a p p e n i n g                                               a n d     n o t     h a p p e n i n g     i s     o n e . ( e )                   0 = N o     c h a n c e     o f     h a p p e n i n g .               H e n c e ,     ( a ) ↔ ( i v ) , ( b ) ↔ ( v ) , ( c ) ↔ ( i ) , ( d ) ↔ ( i i i ) , ( e ) ↔ ( i i )

2. Match the following

(a) If E1 and E2 are the two mutually                  (i) E1 ∩ E2 = E1 exclusive events

(b) If E1 and E2 are the mutually                         (ii) (E1 – E2) ∪ (E1 ∩ E2) = E1

      exclusive and exhaustive events

(c) If E1 and E2 have common                            (iii) E1 ∩ E2 = φ, E1 ∪ E2 = S

     outcomes, then

(d) If E1 and E2 are two events                           (iv) E1 ∩ E2 = φ

      such that E1 ⊂ E2

Sol. ( a ) I f     E 1     a n d     E 2     a r e     m u t u a l l y     e x c l u s i v e     e v e n t s ,     t h e n     E 1 ∩ E 2 = ϕ . ( b ) I f     E 1     a n d     E 2     a r e     m u t u a l l y     e x c l u s i v e     a n d     e x h a u s t i v e     e v e n t s ,     t h e n             E 1 ∩ E 2 = ϕ     a n d     E 1 ∪ E 2 = S . ( c ) I f     E 1     a n d     E 2     h a v e     c o m m o n     o u t c o m e s ,     t h e n     ( E 1 − E 2 ) ∪ ( E 1 ∩ E 2 ) = E 1 ( d ) I f     E 1     a n d     E 2     a r e     t w o     e v e n t s     s u c h     t h a t     E 1 ⊂ E 2         ⇒ E 1 ∩ E 2 = E 1               H e n c e ,     ( a ) ↔ ( i v ) , ( b ) ↔ ( i i i ) , ( c ) ↔ ( i i ) , ( d ) ↔ ( i )
Q&A Icon
Commonly asked questions
Q:  

Match the following

(a) If E1 and E2 are the two mutually                  (i) E1 ∩ E2 = E1 exclusive events

(b) If E1 and E2 are the mutually                         (ii) (E1 – E2) ∪ (E1 ∩ E2) = E1

      exclusive and exhaustive events

(c) If E1 and E2 have common                            (iii) E1 ∩ E2 = φ, E1 ∪ E2 = S

     outcomes, then

(d) If E1 and E2 are two events                           (iv) E1 ∩ E2 = φ

      such that E1 ⊂ E2

Read more
A: 

This is a matching answer type question as classified in NCERT Exemplar

( a ) I f     E 1     a n d     E 2     a r e     m u t u a l l y     e x c l u s i v e     e v e n t s ,     t h e n     E 1 ∩ E 2 = ? . ( b ) I f     E 1     a n d     E 2     a r e     m u t u a l l y     e x c l u s i v e     a n d     e x h a u s t i v e     e v e n t s ,     t h e n             E 1 ∩ E 2 = ?     a n d     E 1 ∪ E 2 = S . ( c ) I f     E 1     a n d     E 2     h a v e     c o m m o n     o u t c o m e s ,     t h e n     ( E 1 − E 2 ) ∪ ( E 1 ∩ E 2 ) = E 1 ( d ) I f     E 1     a n d     E 2     a r e     t w o     e v e n t s     s u c h     t h a t     E 1 ⊂ E 2         ⇒ E 1 ∩ E 2 = E 1               H e n c e ,     ( a ) ↔ ( i v ) , ( b ) ↔ ( i i i ) , ( c ) ↔ ( i i ) , ( d ) ↔ ( i )

Q:  

Match the proposed probability under Column C1 with the appropriate written description under column C2 :

C1                                                                    C2

Probability                                                       Written Description

(a) 0.95                                                             (i) An incorrect assignment

(b) 0.02                                                             (ii) No chance of happening

(c) – 0.3                                                            (iii) As much chance of happening as not.

(d) 0.5                                                               (iv) Very likely to happen

(e) 0                                                                  (v) Very little chance of happening

Read more
A: 

This is a matching answer type question as classified in NCERT Exemplar

( a ) 0 . 9 5 = V e r y     l i k e l y     t o     h a p p e n ,     s o     i t     i s     c l o s e     t o     1 . ( b ) 0 . 0 2 = V e r y     l i t t l e     c h a n c e     o f     h a p p e n i n g     a s     t h e     p r o b a b i l i t y     i s     v e r y     l o w . ( c ) − 0 . 3 = A n     i n c o r r e c t     a s s i g n m e n t     b e c a u s e     p r o b a b i l i t y     i s     n e v e r     n e g a t i v e . ( d )         0 . 5 = A s     m u c h     c h a n c e     o f     h a p p e n i n g     a s     n o t     b e c a u s e     s u m     o f     c h a n c e s     o f     h a p p e n i n g                                               a n d     n o t     h a p p e n i n g     i s     o n e . ( e )                   0 = N o     c h a n c e     o f     h a p p e n i n g .               H e n c e ,     ( a ) ↔ ( i v ) , ( b ) ↔ ( v ) , ( c ) ↔ ( i ) , ( d ) ↔ ( i i i ) , ( e ) ↔ ( i i )

qna

Maths NCERT Exemplar Solutions Class 11th Chapter Sixteen Exam

Student Forum

chatAnything you would want to ask experts?
Write here...